# This work is supported by the ML4Seismic Industry Partners at Georgia Tech

### v1

@@ II-D Inferential Behavior
In seismic fault delineation, these inferential tendencies are amplified by the thin, curvilinear nature of faults and the high noise intrinsic to seismic volumes [63]. U-Net’s symmetric encoder–decoder and skip connections excel at preserving local detail, producing crisp fault traces when the signal‐to‐noise ratio is high [31], but its reliance on local convolutions and fixed strides can fragment continuous faults under heteroskedastic noise [47]. DeepLab’s atrous convolutions and Atrous Spatial Pyramid Pooling (ASPP) module gather multi‐scale context, yielding smoother, globally coherent fault masks [29]; yet the dilation patterns can inadvertently merge adjacent non‐fault discontinuities, introducing false positives along stratigraphic horizons [29]. SegFormer’s transformer‐based encoder captures long‐range dependencies and adapts to complex fault geometries, enhancing continuity across sections [39]—though its patch‐based attention can produce coarser boundaries if patch size is not carefully chosen [39]. These model‐specific inferential biasesdrive distinct behaviors in fault delineation and inform the choice of architectures and post‐processing strategies for robust seismic interpretation.
@@ II-E Performance Evaluation
Segmentation evaluation seeks to quantify how closely a predicted mask $P$ approximates the ground-truth mask $G$ as subsets of the image domain. Evaluation metrics are categorized in the following two ways: (i) region based measures, which treat masks as sets and quantify volumetric overlap, and (ii) distance-based measures, which treat masks as shapes and quantify geometric deviations along their contours.
Region-based metrics include the Jaccard index (also known as intersection-over-union):
$\mathrm{Jaccard}(P,G)=\frac{\lvert P\cap G\rvert}{\lvert P\cup G\rvert},$ (5)
and the Dice coefficient:
$\mathrm{Dice}(P,G)\;=\;\frac{2\,\lvert P\cap G\rvert}{\lvert P\rvert+\lvert G\rvert},$ (6)
which captures the proportion of shared pixels but may understate errors on thin or tortuous structures.
Distance-based metrics such as Hausdorff offer tolerance against spatial shifts, but they are sensitive to noisy outliers [72]. To mitigate this issue, the modified Hausdorff was proposed by Dubuisson and Jain [73] that considers the average distance to the nearest neighbor, rather than the maximum distance as in the traditional formulation. In [74], the authors utilize the Bidirectional Chamfer Distance (BCD) metric, which measures the minimum distance from each pixel in both directions. Although region based metrics such as the Dice coefficient are useful for assessing shape quality, from a geophysicist’s perspective, evaluation should measure the structural similarity between lines rather than discrete points. The modified Hausdorff distance $D_{h}$ and the BCD $D_{c}$ between two non-empty sets $P$ and $G$ are defined respectively as:
$D_{h}\bigl{(}P,G\bigr{)}$ $=\max\Biggl{(}\frac{1}{N_{p}}\sum_{p\in P}\min_{g\in G}\|p-g\|,$
$\hskip 41.00012pt\frac{1}{N_{g}}\sum_{g\in G}\min_{p\in P}\|g-p\|\Biggr{)}$ (7)
$D_{c}(P,G)$ $=\frac{1}{N_{p}}\sum_{p\in G}\min_{g\in G}\|p-g\|^{2}+\frac{1}{N_{g}}\sum_{g\in G}\min_{p\in P}\|g-p\|^{2}$ (8)
where:
[…]
  • $\min_{g\in G}\|p-g\|$ is the distance from point $p\in P$ to its nearest neighbor in $G$,
  • $\sum_{p\in P}$ and $\sum_{g\in G}$ are summations over all points in sets $P$ and $G$, respectively,
  • ${N_{p}}$ and ${N_{g}}$ are the number of elements in the sets $P$ and $G$, respectively.
To illustrate the complementary nature of these metrics, we present a toy example in Fig. 4. The Dice coefficient measures spatial overlap by comparing each predicted pixel against the ground truth, yielding a single overlap score. In contrast, both the bidirectional Chamfer distance (BCD) and the modified Hausdorff distance assess boundary discrepancies: BCD sums the forward and reverse distances, whereas the modified Hausdorff takes the larger of the two. Because BCD and the modified Hausdorff yield nearly identical trends, our figures plot either one of them alongside the Dice coefficient.
Figure 4: Differences between the considered metrics.
@@ III Fault Delineation Datasets
Besides the lack of public availability, the different characteristics of the datasets pose challenges to the development of generalizable ML models. Specifically, LANDMASS contains image-level fault labels that cannot be used to numerically evaluate the delineation of pixel-wise faults. GSB contains pixel-wise fault labels annotated on only $5$ crossline sections, which limits the evaluation scalability on large test data. Additionally, it is challenging to achieve generalization using only a small number of labels for finetuning [50]. In contrast, Thebe provides a large amount of pixel-level geophysicist labels across $1803$ crossline sections. CRACKS provides fault labels of varying quality collected from a group of interpreters, including a geophysicist expert, across $400$ inline sections.
CRACKS is an open-source dataset with diverse faults delineated across $400$ inline sections of the Netherlands F3 Block [86, 90]. The authors in [10] open-sourced a fully-annotated 3D geological volume of the Netherlands F3 Block for training different models and comparing the performance with objective metrics. Thus, this volume is one of the most extensively studied geographical zones for developing ML-assisted seismic interpretation frameworks [2, 48, 50, 22, 91, 49, 20, 52, 12, 13, 92, 93, 94, 95, 53, 96, 14, 97, 98, 82, 81, 99, 100, 101, 10, 44, 45, 102, 46, 103, 104, 105, 106]. The diverse fault features in the F3 block, including major versus minor faults and varying orientations, make it an excellent seismic dataset to train and evaluate fault delineation models [107, 108]. However, the annotations in [10] do not provide pixel-wise fault labels. Thus, CRACKS open-sources fully hand-labeled fault annotations by a group of $32$ interpreters with varying degrees of expertise and a domain expert geophysicist. This dataset not only establishes a standardized benchmark for objective comparison but also can be used to investigate the impact of multiple sets of labels with varying quality. Three sets of fault annotations are used to investigate the impact of training labels with varying quality, including expert labels and two more sets of lower quality labels from two other annotators. Fig. 5(b) shows an example inline section with expert fault labels from CRACKS. CRACKS and FaultSeg3D share geological similarity in the seismic sections in addition to the similar label density.
@@
Figure 5: Visuals from three datasets: (5(a)) synthesized faults in FaultSeg3D, (5(b)) expert labels in CRACKS, and (5(c)) expert labels in Thebe.
@@ IV Experimental Setup
The workflow of ML-assisted fault delineation involves multiple choices including ($i$) the decision to fine-tune or not, ($ii$) the selection of the datasets for pre-training and fine-tuning, ($iii$) the selection of different models, ($iv$) the development of pre-processing and post-processing strategies, and ($v$) the standardization of the evaluation protocols. We depict our systematic benchmarking pipeline in Fig. 6, which we use to holistically compare the impact of every choice and the various combinations of such choices. We provide details on each component of our pipeline in this section.
@@ IV-A1 Standardizing Fault Annotations using Morphological Operations
There exist significant differences between the thickness of fault annotations in CRACKS, Thebe, and FaultSeg3D. We show three labeled patches of the same size from Thebe, CRACKS, and FaultSeg3D in Fig. 7(a). The faults manually delineated by interpreters in Thebe and CRACKS are thicker than the synthesized faults in FaultSeg3D. The inconsistent thickness of fault annotations often influences model optimization, which typically uses pixel-based loss functions. However, existing workflows neglect this fault annotation inconsistency [109, 88, 84, 85]. The synthesized faults in FaultSeg3D precisely delineate the thin fractures in the seismic volumes. We therefore standardize the fault thickness across datasets by processing the manually delineated faults in Thebe and CRACKS to have the same thickness of the faults in FaultSeg3D. Specifically, we apply a simple sequence of morphological operations: we first skeletonize the raw fault annotations and then apply dilation with a rank-3 structural element to uniformize fault thickness and close any small gaps, as shown in Fig. 7(b). This process is denoted by the Preprocess block in Fig. 6. We use standard Python libraries scikit-image and scipy for these morphological operations.
@@ IV-C2 Performance Metrics
As mentioned in Section II-E, pixel-based and distance-based metrics capture different aspects of prediction quality, and each can fail to fully reflect the structural accuracy of the predicted faults, making the evaluation of fault delineation methods an inherently challenging task. Our benchmark study thus adopts a holistic approach to assess prediction quality by considering a combination of multiple metrics alongside subjective inspection. In this study we choose to report one pixel-based metric, Dice Coefficient (defined in Eq. (6)) and two distance-based metrics: BCD and the modified Hausdorff (defined in Eq. 8, and Eq. 7, respectively), whose difference we showcased in Fig. 4. These metrics have not only been extensively used in the seismic literature for model performance assessment [1, 25, 73], but allow for two different evaluation axes: pixel overlap-based and structure-based.
@@ V Results
Benchmarking results are analyzed across three different thematic axes: (1) generalizability and transferability, (2) training dynamics, and (3) metric evaluation.
@@ V-A1 Dataset Alignment and Transfer Trends
Figs. 910 and 11 provide a macroscopic view of the model behaviors across the different training setups. On CRACKS (Fig. 9), the top-performing configurations are those pretrained on FaultSeg3D data and fine-tuned on CRACKS. The observation indicates strong geophysical commonalities between FaultSeg3D and CRACKS data, and supports the utility of FaultSeg3D data as a viable pretraining source when real annotations are limited. The fill-color distribution in Fig. 9 also establishes a hierarchy of effective fine-tuning datasets for CRACKS: CRACKS ¿ FaultSeg3D ¿ Thebe. The poor finetuning performance on Thebe indicates that it is more distributionally distant from CRACKS than FaultSeg3D data.
Furthermore, when tested on FaultSeg3D data (Fig. 10), models pretrained on CRACKS again outperform those trained from scratch, indicating that the aforementioned alignment is reciprocal. However, for Thebe (Fig. 11), top-performing models are those trained from scratch on Thebe itself. Transferring from either FaultSeg3D or CRACKS results in performance degradation, suggesting that Thebe resides in a distinct feature space.
Figure 9: Test on CRACKS dataset
Figure 10: Test on FaultSeg3D dataset
Figure 11: Test on Thebe dataset
@@ V-A2 Domain Shift and Joint Training
Domain shift is a pressing challenge in seismic ML, particularly when deploying models across surveys with different geological properties. When overlooked, this can lead to brittle models that perform well on training data but fail on new volumes, or models that catastrophically forget features from their original data after finetuning. An illustration of this catastrophic forgetting occurs when models pretrained on CRACKS are fine-tuned on Thebe: despite reasonable performance on Thebe, these models experience dramatic performance degradation when re-evaluated on their original domain. In the case of unet++, for instance, the Dice score on CRACKS drops from 0.34 to 0.12, while the BCD increases twofold, indicating a complete erasure of useful representations, and clear case of catastrophic forgetting. A statistical analysis shows that Thebe has a significantly lower standard deviation (0.124) compared to CRACKS (1.149) and FaultSeg3D (1.052). The contrast and intensity variations in Thebe are lower, while CRACKS and FaultSeg3D have diverse intensity distributions, as shown in Fig. 5. An explanation for this catastrophic forgetting phenomenon is that models trained on Thebe learn from a constrained input range and struggle when tested on datasets with richer intensity distributions. These results suggest that other normalization techniques before training could help reduce these distributional mismatches.
As an additional baseline, experiments on joint training on all possible combinations of datasets in a single training round are conducted using the unet model with a ResNet50 backbone. The results for all unet experiments, including these joint training configurations, are shown in Table I. The results further showcase that features learned from CRACKS and FaultSeg3D data align together, but these learned features are not easily transferable to Thebe. On the other hand, Thebe acts as a regularizer, hurting performance on seen datasets but providing modest generalization to unseen ones.
@@
FaultSeg3D + Cracks 0.662$\pm$0.190 6.143$\pm$20.421 5.282$\pm$18.295 0.333$\pm$0.048 15.436$\pm$5.254 11.597$\pm$4.836 0.001$\pm$0.001 659.708$\pm$74.985 430.887$\pm$64.880
FaultSeg3D + Thebe 0.060$\pm$0.031 29.834$\pm$20.870 20.151$\pm$19.805 0.167$\pm$0.033 29.717$\pm$8.512 26.136$\pm$8.187 0.211$\pm$0.024 52.952$\pm$10.802 43.348$\pm$10.516
Cracks + Thebe 0.038$\pm$0.3375 34.183$\pm$28.569 22.751$\pm$23.675 0.32$\pm$0.067 13.797$\pm$6.2072 8.2964$\pm$4.64058 0.112$\pm$0.03 112.354$\pm$38.606 60.198$\pm$40.66
All 0.337$\pm$0.035 28.569$\pm$13.036 23.674$\pm$11.087 0.306$\pm$0.0743 17.7453$\pm$21.851 12.005$\pm$19.724 0.144$\pm$0.029 75.877$\pm$37.191 59.039$\pm$19.8
@@
Cracks $\rightarrow$ FaultSeg3D 0.667$\pm$0.171 4.994$\pm$16.049 3.821$\pm$10.475 0.028$\pm$0.009 73.433$\pm$13.961 45.378$\pm$13.263 0.027$\pm$0.003 165.315$\pm$10.283 99.996$\pm$7.698
Cracks $\rightarrow$ Thebe 0.099$\pm$0.036 21.959$\pm$13.845 15.771$\pm$10.753 0.159$\pm$0.028 26.070$\pm$7.107 20.929$\pm$6.552 0.181$\pm$0.014 53.808$\pm$10.361 42.169$\pm$12.450
FaultSeg3D $\rightarrow$ Cracks 0.082$\pm$0.080 136.293$\pm$151.136 134.095$\pm$151.213 0.375$\pm$0.059 8.806$\pm$3.728 5.577$\pm$2.71 0.019$\pm$0.005 213.695$\pm$20.154 147.213$\pm$12.756
FaultSeg3D $\rightarrow$ Thebe 0.075$\pm$0.082 93.488$\pm$66.848 84.420$\pm$61.439 0.135$\pm$0.023 28.701$\pm$7.022 23.368$\pm$7.111 0.158$\pm$0.014 78.683$\pm$35.415 51.649$\pm$30.973
Thebe $\rightarrow$ Cracks 0.019$\pm$0.027 56.145$\pm$34.593 42.899$\pm$34.785 0.317$\pm$0.047 10.896$\pm$3.127 6.656$\pm$2.500 0.019$\pm$0.005 161.131$\pm$14.429 88.344$\pm$9.217
Thebe $\rightarrow$ FaultSeg3D 0.526$\pm$0.203 9.319$\pm$22.281 8.182$\pm$21.079 0.053$\pm$0.012 38.141$\pm$7.286 34.319$\pm$7.233 0.037$\pm$0.003 134.522$\pm$7.354 122.936$\pm$7.078

### v2

@@ II-D Inferential Behavior
In seismic fault delineation, these inferential tendencies are amplified by the thin, curvilinear nature of faults and the high noise intrinsic to seismic volumes [77]. U-Net’s symmetric encoder–decoder and skip connections excel at preserving local detail, producing crisp fault traces when the signal‐to‐noise ratio is high [45], but its reliance on local convolutions and fixed strides can fragment continuous faults under heteroskedastic noise [61]. DeepLab’s [43] atrous convolutions and Atrous Spatial Pyramid Pooling (ASPP) module gather multi‐scale context, yielding smoother, globally coherent fault masks; yet the dilation patterns can inadvertently merge adjacent non‐fault discontinuities, introducing false positives along stratigraphic horizons. SegFormer’s transformer‐based encoder captures long‐range dependencies and adapts to complex fault geometries enhancing continuity across sections, though its patch‐based attention can produce coarser boundaries if patch size is not carefully chosen [53].
Furthermore, recent works in both seismic [78, 79] and general DL [89, 90] domains emphasize that inferential biases are not merely architectural, but also dataset-driven: synthetic datasets may thus encourage smoother, well-connected predictions, while field datasets can induce models to replicate jagged, discontinuous, or crossover-heavy structures (we analyze these trends in depth using the fault characteristic metrics described in Section II-F). New approaches attempt to counteract these biases by introducing geological priors, like fault continuity preservation [20] or curvature-based losses [34]. These studies highlight that inferential behavior in seismic fault delineation is shaped jointly by architectural design and the structural biases present in training data, underscoring the importance of evaluation frameworks that go beyond pixel overlap to account for geological plausibility.
@@ II-E Performance Evaluation
Evaluation metrics for fault interpretation can be broadly grouped into two categories: region-based and distance-based metrics. Region-based metrics quantify the overlap between predicted and reference fault regions, while distance-based metrics measure the geometric discrepancy between their boundaries.
Region-based metrics. The Dice coefficient $D$ measures the spatial overlap between prediction $P$ and ground truth $G$
$D\;=\;\frac{2\,\lvert P\cap G\rvert}{\lvert P\rvert+\lvert G\rvert}$ (1)
Distance-based metrics. The modified Hausdorff distance $HD$ captures the largest average boundary deviation:
$HD=\max\Biggl{(}\frac{1}{N_{p}}\sum_{p\in P}\min_{g\in G}d(p,g),\frac{1}{N_{g}}\sum_{g\in G}\min_{p\in P}d(p,g)\Biggr{)}$ (2)
where $d(\cdot,\cdot)$ is the Euclidean distance between two points, and $N_{p}$} and $N_{g}$} are the number of elements in the sets $P$ and $G$, respectively. Additionally, the Bidirectional Chamfer Distance $BCD$ computes the average shortest distance between boundary points in both directions:
$BCD=\frac{1}{N_{p}}\sum_{p\in G}\min_{g\in G}d(p,g)+\frac{1}{N_{g}}\sum_{g\in G}\min_{p\in P}d(p,g)$ (3)
This categorization provides a basis for later discussion in the Results section, where we compare the behaviors of overlap- and distance-focused metrics.
II-F Fault Characteristics Metrics
To characterize fault networks within each dataset, we measure a set of geometric and topological descriptors. These include Length, Curvature, Sinuosity, Segments, and Stepover Density, which correspond to the statistics reported in Table III.
Length ($L$) [91] is the total length of all fault traces, computed as
$L=\sum_{i=1}^{N_{\text{faults}}}\ell_{i},$ (4)
where $\ell_{i}$ denotes the length of the $i$-th fault and $N_{\text{faults}}$ is the number of fault traces.
Curvature ($\kappa$) [92] quantifies the local bending of fault traces. At each arc-length position $s$, curvature is defined as
$\kappa(s)=\left|\frac{d\theta}{ds}\right|,$ (5)
where $\theta$ is the tangent orientation along the fault.
Sinuosity ($S$) [93] measures how tortuous a fault is, defined as the ratio of its length to the straight-line distance between its endpoints:
$S=\frac{L_{\text{trace}}}{D_{\text{endpoints}}}.$ (6)
Segments ($N_{\text{seg}}$) [94] is the total number of discrete fault segments observed in the network.
Stepover Density ($D_{\text{stepover}}$) [95] measures the relative occurrence of stepovers, normalized by fault length:
$D_{\text{stepover}}=\frac{N_{\text{stepover}}}{L},$ (7)
where $N_{\text{stepover}}$ is the total number of stepovers. These descriptors enable dataset-level comparisons of fault geometry and topology.
@@ III Fault Delineation Datasets
Figure 3: Visuals from three datasets: (3(a)) synthesized faults in FaultSeg3D, (3(b)) expert labels in CRACKS, and (3(c)) expert labels in Thebe.
Besides the lack of public availability, the different characteristics of the datasets pose challenges to the development of generalizable DL models. Specifically, LANDMASS contains image-level fault labels that cannot be used to numerically evaluate the delineation of pixel-wise faults. GSB contains pixel-wise fault labels annotated on only $5$ crossline sections, which limits the evaluation scalability on large test data. Additionally, it is challenging to achieve generalization using only a small number of labels for finetuning [64]. In contrast, Thebe provides a large amount of pixel-level geophysicist labels across $1803$ crossline sections. CRACKS provides fault labels of varying quality collected from a group of interpreters, including a geophysicist expert, across $400$ inline sections.
CRACKS is an open-source dataset with diverse faults delineated across $400$ inline sections of the Netherlands F3 Block [107, 111]. The authors in [24] open-sourced a fully-annotated 3D geological volume of the Netherlands F3 Block for training different models and comparing the performance with objective metrics. Thus, this volume is one of the most extensively studied geographical zones for developing DL-assisted seismic interpretation frameworks [16, 62, 64, 36, 112, 63, 34, 66, 26, 27, 113, 114, 115, 116, 67, 117, 28, 118, 119, 103, 102, 120, 121, 122, 24, 58, 59, 123, 60, 124, 125, 126, 127]. The diverse fault features in the F3 block, including major versus minor faults and varying orientations, make it an excellent seismic dataset to train and evaluate fault delineation models [128, 129]. However, the annotations in [24] do not provide pixel-wise fault labels. Thus, CRACKS open-sources fully hand-labeled fault annotations by a group of $32$ interpreters with varying degrees of expertise and a domain expert geophysicist. This dataset not only establishes a standardized benchmark for objective comparison but also can be used to investigate the impact of multiple sets of labels with varying quality. Three sets of fault annotations are used to investigate the impact of training labels with varying quality, including expert labels and two more sets of lower quality labels from two other annotators. Fig. 3(b) shows an example inline section with expert fault labels from CRACKS. CRACKS and FaultSeg3D share geological similarity in the seismic sections in addition to the similar label density.
@@ IV Experimental Setup
The workflow of DL-assisted fault delineation involves multiple choices including ($i$) the decision to fine-tune or not, ($ii$) the selection of the datasets for pre-training and fine-tuning, (iiiiii) the selection of different models, ($iv$) the development of pre-processing and post-processing strategies, and ($v$) the standardization of the evaluation protocols. We summarize these steps in Fig. 4, where the pipeline is organized chronologically in a top-to-bottom fashion to reflect the order in which decisions are made in practice. This systematic layout enables us to holistically compare the effect of each component and their combinations. We provide details on each stage of the pipeline in the remainder of this section.
@@ IV-A1 Standardizing Fault Annotations using Morphological Operations
There exist significant differences between the thickness of fault annotations in CRACKS, Thebe, and FaultSeg3D. As shown in Fig.5(a), the manually delineated faults in Thebe and CRACKS are considerably thicker than the synthesized faults in FaultSeg3D. Such inconsistency can systematically bias model training and evaluation, especially under pixel-wise loss functions such as the Dice loss. In particular, Dice loss computes similarity between predicted and reference masks by balancing overlap against the number of positive pixels in each mask. Thicker annotations artificially increase the proportion of positive pixels, making the loss less sensitive to small spatial deviations and allowing models to achieve high scores without precisely localizing the fault centerline. Conversely, thinner annotations lead to a stronger penalization of misalignments, requiring sharper localization for similar Dice scores.
Without standardization, these annotation thickness differences could confound cross-dataset comparisons in our benchmark: a model evaluated on thicker annotations may appear more accurate than one evaluated on thinner ones, even if their true localization capability differs. To mitigate this bias, we standardize fault thickness across datasets by processing the manually delineated faults in Thebe and CRACKS to match the thinner style of FaultSeg3D. Specifically, we skeletonize the raw annotations and then apply dilation with a rank-3 structural element to produce uniform thickness and close small gaps, as illustrated in Fig.5(b). This preprocessing, shown in the Preprocess block in Fig.4, is implemented with scikit-image and scipy.
@@ IV-C2 Performance Metrics
As mentioned in Section II-E, pixel-based and distance-based metrics capture different aspects of prediction quality, and each can fail to fully reflect the structural accuracy of the predicted faults, making the evaluation of fault delineation methods an inherently challenging task. Our benchmark study thus adopts a holistic approach to assess prediction quality by considering a combination of multiple metrics alongside subjective inspection. In this study we choose to report one pixel-based metric, Dice Coefficient (defined in Eq. (1)) and two distance-based metrics: BCD and the modified Hausdorff (defined in Eq. 3, and Eq. 2, respectively). These metrics have not only been extensively used in the seismic literature for model performance assessment [15, 39, 137], but allow for two different evaluation axes: pixel overlap-based and structure-based.
IV-C3 Fault Characteristic Comparisons
While the individual metrics described in Section II-F quantify dataset characteristics in isolation, comparative metrics evaluate the similarity between predicted and ground-truth faults. We describe below the metrics we use in Table IV to extend the individual descriptors into differences, ratios, and statistical comparisons.
Strike Similarity (StrikeSim) measures the similarity of orientation distributions between predicted ($P(\theta)$) and ground truth ($G(\theta)$) faults using cosine similarity:
$\text{StrikeSim}=\frac{\sum_{\theta}P(\theta)G(\theta)}{\sqrt{\sum_{\theta}P(\theta)^{2}}\,\sqrt{\sum_{\theta}G(\theta)^{2}}}.$ (8)
Curvature Metrics compare curvature distributions across predictions and ground truth. These include:
$\Delta\kappa$ $=\big{|}\overline{\kappa}_{\text{pred}}-\overline{\kappa}_{\text{gt}}\big{|},$ (9)
$\text{RMSE}_{\kappa}$ $=\sqrt{\frac{1}{N}\sum_{i=1}^{N}\left(\kappa_{i}^{\text{pred}}-\kappa_{i}^{\text{gt}}\right)^{2}},$ (10)
$\text{Corr}_{\kappa}$ $=\frac{\text{Cov}\!\left(\kappa^{\text{pred}},\kappa^{\text{gt}}\right)}{\sigma_{\kappa^{\text{pred}}}\sigma_{\kappa^{\text{gt}}}}.$ (11)
Sinuosity Metrics evaluate differences in tortuosity:
$\Delta S$ $=\overline{S}_{\text{pred}}-\overline{S}_{\text{gt}},$ (12)
$R_{S}$ $=\frac{\overline{S}_{\text{pred}}}{\overline{S}_{\text{gt}}}.$ (13)
Length Metrics evaluate differences in fault trace length:
$\Delta L$ $=L_{\text{pred}}-L_{\text{gt}},$ (14)
$R_{L}$ $=\frac{L_{\text{pred}}}{L_{\text{gt}}}.$ (15)
Segment Metrics compare the number of interpreted fault segments:
$\Delta N_{\text{seg}}$ $=N_{\text{seg}}^{\text{pred}}-N_{\text{seg}}^{\text{gt}},$ (16)
$R_{N_{\text{seg}}}$ $=\frac{N_{\text{seg}}^{\text{pred}}}{N_{\text{seg}}^{\text{gt}}}.$ (17)
Stepover Metrics compare the relative frequency of stepovers:
$\Delta D_{\text{stepover}}$ $=D_{\text{stepover}}^{\text{pred}}-D_{\text{stepover}}^{\text{gt}},$ (18)
$R_{D_{\text{stepover}}}$ $=\frac{D_{\text{stepover}}^{\text{pred}}}{D_{\text{stepover}}^{\text{gt}}}.$ (19)
By combining individual metrics with their comparative counterparts, we assess not only whether faults are detected, but also whether their structural and geometric properties are faithfully reproduced.
@@ V Results
The results of our benchmarking experiments are analyzed across three different thematic axes: (1) generalizability and transferability, (2) training dynamics, and (3) metric evaluation.
@@ V-A1 Dataset Alignment and Transfer Trends
Fig. 7 provides a macroscopic view of the model behaviors across the different training setups. On CRACKS (Fig. 7(a)), the top-performing configurations are those pretrained on FaultSeg3D data and fine-tuned on CRACKS. The observation indicates strong geophysical commonalities between FaultSeg3D and CRACKS data, and supports the utility of FaultSeg3D data as a viable pretraining source when real annotations are limited. The fill-color distribution in Fig. 7(a) also establishes a hierarchy of effective fine-tuning datasets for CRACKS: CRACKS ¿ FaultSeg3D ¿ Thebe. The poor finetuning performance on Thebe indicates that it is more distributionally distant from CRACKS than FaultSeg3D data.
Furthermore, when tested on FaultSeg3D data (Fig. 7(b)), models pretrained on CRACKS again outperform those trained from scratch, indicating that the aforementioned alignment is reciprocal. However, for Thebe (Fig. 7(c)), top-performing models are those trained from scratch on Thebe itself. Transferring from either FaultSeg3D or CRACKS results in performance degradation, suggesting that Thebe resides in a distinct feature space.
(a) Test on CRACKS dataset
(b) Test on FaultSeg3D dataset
(c) Test on Thebe dataset
Figure 7: Top: All of out models across our different pretraining and finetuning setups. Bottom: Best performing models for each dataset.
@@ V-A2 Domain Shift and Joint Training
Domain shift is a pressing challenge in seismic DL, particularly when deploying models across surveys with different geological properties. When overlooked, this can lead to brittle models that perform well on training data but fail on new volumes, or models that catastrophically forget features from their original data after finetuning. An illustration of this catastrophic forgetting occurs when models pretrained on CRACKS are fine-tuned on Thebe: despite reasonable performance on Thebe, these models experience dramatic performance degradation when re-evaluated on their original domain. In the case of unet++, for instance, the Dice score on CRACKS drops from 0.34 to 0.12, while the BCD increases twofold, indicating a complete erasure of useful representations, and clear case of catastrophic forgetting. A statistical analysis shows that Thebe has a significantly lower standard deviation (0.124) compared to CRACKS (1.149) and FaultSeg3D (1.052). The contrast and intensity variations in Thebe are lower, while CRACKS and FaultSeg3D have diverse intensity distributions, as shown in Fig. 3. An explanation for this catastrophic forgetting phenomenon is that models trained on Thebe learn from a constrained input range and struggle when tested on datasets with richer intensity distributions. These results suggest that other normalization techniques before training could help reduce these distributional mismatches.
As an additional baseline, experiments on joint training on all possible combinations of datasets in a single training round are conducted using the unet model with a ResNet50 backbone. The results for all unet experiments, including these joint training configurations, are shown in Table I. The results further showcase that features learned from CRACKS and FaultSeg3D data align together, but these learned features are not easily transferable to Thebe. On the other hand, Thebe may act as a regularizer when paired with a large dataset like FaultSeg3D, hurting performance on seen datasets (compared to training them individually) but providing modest generalization to unseen ones: performance on the unseen dataset for this setting (i.e. CRACKS) does not drop as badly as that of the unseen dataset for other joint settings (e.g. FaultSeg3D + CRACKS).
@@
[…]
@@
[…]